### v1

@@ 6 Improved Tree Cut-Sparsifier
We begin by re-introducing the relevant notation from Section 5. As per Corollary 5.3, we fix a cut $(B\subseteq V,W=V\setminus B)$ in $G$ and a valid demand state $P$ on $V$, which is $1$-respected by $T$. We work in the subdivision graph $G^{\prime}=(V^{\prime}\cup X_{E},E^{\prime})$. We define for a cluster $S\subseteq V$, the corresponding cluster in the subdivision graph as $S^{\prime}\mathrel{\mathop{:}}=S\cup\{x_{e}\mid e\cap S\geq 1\}$ and let $G^{\prime}[S^{\prime}]$ be the subdivision graph restricted to the set $S^{\prime}$. We let $(B^{\prime}=B\cup\{x_{e}\mid e\cap B\geq 1\},W^{\prime}=V^{\prime}\setminus B^{\prime})$ be the cut corresponding to $(B,W)$ in $G^{\prime}$. We will charge the edges $E(B^{\prime},W^{\prime})$ in $G^{\prime}$ to cover the demand $\operatorname{dem}_{P^{\prime}}(B^{\prime},W^{\prime})$, where $P^{\prime}$ will be the appropriate demand state on $G^{\prime}$.
Like in Section 5, we convert the valid demand state $P$ to demand states $P_{\{v\}}^{\prime}$ for each vertex $v\in V$, by distributing the demand vector of $v$ to the split nodes neighboring $v$ evenly. These demand states satisfy Invariant 5.4 with $\alpha=1$. If we denote the sum of these demand states as $P_{\mathcal{Q}_{0}}$, Claim 5.8 shows that $\operatorname{dem}_{P}(B,W)\leq\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(B^{\prime},W^{\prime})+\operatorname{cap}_{G}(B,W)$.
@@
Let $P^{(1)}_{S^{\prime}_{i,j}}\mathrel{\mathop{:}}=P_{S^{\prime}_{i,j}}$. We apply the first guarantee of Theorem 6.2 on each of the clusters $S_{i,j}$ ($1\leq i\leq t,1\leq j\leq t_{i}$), to distribute the demand states $P^{(1)}_{S^{\prime}_{i,j}}$ uniformly on the boundary edges of the cluster. Formally, define the average demand vector $N_{i,j}\mathrel{\mathop{:}}=\frac{1}{|X_{B_{i,j}}|}\sum_{v\in X_{B_{i,j}}}{P^{(1)}_{S^{\prime}_{i,j}}(v,\cdot)}$, where $B_{i,j}=E(S_{i,j},V\setminus S_{i,j})$ are the boundary edges of the cluster $S_{i,j}$. Consider the demand state $P^{(2)}_{S^{\prime}_{i,j}}$ which is supported on $X_{B_{i,j}}$, where each demand vector is $N_{i,j}$. By construction, $P^{(2)}_{S^{\prime}_{i,j}}$ satisfies Invariant 5.4 for $S_{i,j}$ with the same parameter $\alpha$ as $P^{(1)}_{S^{\prime}_{i,j}}$. The updated demand state of $P^{(1)}_{S^{\prime}_{i,j}}$ with respect to the all-to-all flow problem on $X_{B_{i,j}}$ scaled by $\alpha$ is $P^{(2)}_{S^{\prime}_{i,j}}$ (see Remark 3.20). Using the first guarantee of Theorem 6.2, and by Facts 3.19 and 2.7, the valid demand state $P^{(1)}_{S^{\prime}_{i,j}}-P^{(2)}_{S^{\prime}_{i,j}}$ is $\Omega\left(\frac{1}{\alpha\cdot f(S_{i,j})\cdot\log n}\right)$-respected by $G^{\prime}[S^{\prime}_{i,j}]$. Therefore, by charging $O(\alpha\cdot f(S_{i,j})\cdot\log n)$ to each edge of $E(B^{\prime},W^{\prime})\cap G^{\prime}[S_{i,j}^{\prime}]$, we are able to cover the difference in demands $\operatorname{dem}_{P^{(1)}_{S^{\prime}_{i,j}}}(B^{\prime},W^{\prime})-\operatorname{dem}_{P^{(2)}_{S^{\prime}_{i,j}}}(B^{\prime},W^{\prime})$.
@@
The update by the all-to-all flow problem does not change the total demand (i.e., $P^{(1)}_{S^{\prime}_{i,j}}-P^{(2)}_{S^{\prime}_{i,j}}$ is a valid demand state), so we get
@@
Using the claim, we have that the total load $P^{(2)}_{S^{\prime}_{i,j}}$ puts on $X_{B_{i,j}}$ is at most $\operatorname{cap}_{G}(B_{i,j})$, and because the demand vectors are the same, we get that the load on each vertex is at most $1$.
@@
Next, we use the second guarantee from Theorem 6.2. This allows us to route in $G^{\prime}[S^{\prime}_{i}]$, one unit of mass from each $x_{f}\in X_{F_{i}}$ to $X_{B_{i}}$, such that each vertex $x_{b}\in X_{B_{i}}$ receives at most $O(1)$ units of mass. Scaling this flow by $2$ gives that we can route two units from each inter-cluster edge to the boundary. This flow can be routed with congestion $O(\log n)$. Hence, by Fact 2.7, we can replace $P^{(2)}_{S^{\prime}_{i}}$ with the updated demand state obtained after this routing, denoted $P^{(3)}_{S^{\prime}_{i}}$, while incurring a charge of $O(\log n)$ on the edges $E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}_{i}]$. Consequently, $P^{(3)}_{S^{\prime}_{i}}$ satisfies Invariant 5.4 for $S_{i}$ with $\alpha=O(1)$, as the demand state is supported on nodes of $X_{B_{i}}$ and the load on each node is constant.
Denote $P^{(3)}_{S^{\prime}}\mathrel{\mathop{:}}=\sum_{i=1}^{t}{P^{(3)}_{S^{\prime}_{i}}}$. We now use Theorem 6.1 on the merge-phase partition $\{S_{1},\ldots,S_{t}\}$. It allows us to replace $P^{(3)}_{S^{\prime}}$ with the updated $P^{(4)}_{S^{\prime}}$ which satisfies Invariant 5.4 for $S$ with $\alpha=O(1)$, while charging each edge of $E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$ at most $O(\log n)$ to cover the difference in demands.
This $P^{\mathrm{after}}_{S^{\prime}}\mathrel{\mathop{:}}=P^{(4)}_{S^{\prime}}$ is the demand state required by 6.3, and the total charge on each edge of $E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$ is dominated by $O(\alpha\cdot f(S_{i,j})\cdot\log n)$. ∎
Using Theorem 6.3, Theorem 2.8 follows exactly like in Section 5. We give the full details for completeness.
@@
Denote by $\mathcal{Q}$ the set of current “active” clusters, such that initially $\mathcal{Q}_{0}=\{\{v\}\mid v\in V\}$. Then in each iteration we remove $\{S_{i,j}\mid 1\leq i\leq t,1\leq j\leq t_{i}\}$ from $\mathcal{Q}$ and add $S$ to $\mathcal{Q}$. Let $P_{\mathcal{Q}}$ be the sum of all demand states $P_{S^{\prime}}$, for $S\in\mathcal{Q}$. Using Claim 5.7, we have that $P_{\mathcal{Q}}$ is always a valid demand state on $V^{\prime}$. Claim 5.8 shows that $\operatorname{dem}_{P}(B,W)\leq\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(B^{\prime},W^{\prime})+\operatorname{cap}_{G}(B,W)$.
@@
Consider a replacement of $\mathcal{Q}$ by $\mathcal{Q}^{\prime}=\mathcal{Q}\cup\{S\}\setminus\{S_{i,j}\mid 1\leq i\leq t,1\leq j\leq t_{i}\}$ when we apply Theorem 6.3. Recall that $P^{\mathrm{before}}_{S^{\prime}}=\sum_{i=1}^{t}\sum_{j=1}^{t_{i}}P_{S^{\prime}_{i,j}}$ is replaced by $P^{\mathrm{after}}_{S^{\prime}}$. Let $P^{\mathrm{before}}_{\mathcal{Q}^{\prime}}\mathrel{\mathop{:}}=P_{\mathcal{Q}}$ be the sum of active demand states before the change and $P^{\mathrm{after}}_{\mathcal{Q}^{\prime}}\mathrel{\mathop{:}}=P_{\mathcal{Q}}-P^{\mathrm{before}}_{S^{\prime}}+P^{\mathrm{after}}_{S^{\prime}}$ be the sum after the change. We have $\operatorname{dem}_{\left(P^{\mathrm{before}}_{\mathcal{Q}^{\prime}}-P^{\mathrm{after}}_{\mathcal{Q}^{\prime}}\right)}(B^{\prime},W^{\prime})=\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(B^{\prime},W^{\prime})$. This means that after accumulating all the charges in the applications of Theorem 5.5 on all clusters $S\in T$, the total charge satisfies
$\sum_{e\in E(B^{\prime},W^{\prime})}{\mathrm{charge}(e)}$ $=\sum_{e\in E(B^{\prime},W^{\prime})}\sum_{S\in T}{\mathrm{charge}_{S}(e)}=\sum_{S\in T}\sum_{e\in E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]}{\mathrm{charge}_{S}(e)}$
@@
$\geq\sum_{S\in T}{\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(B^{\prime},W^{\prime})}\geq\operatorname{dem}_{\left(P_{\mathcal{Q}_{0}}-P_{V^{\prime}}\right)}(B^{\prime},W^{\prime}).$
@@
where $\mathrm{charge}_{S}$ is the charge incurred when applying the theorem on the cluster $S$, and the last inequality uses Fact 3.17. Next, note that by Invariant 5.4(1), because $P_{V^{\prime}}$ satisfies the invariant for $V$, it holds that $P_{V^{\prime}}=0$. Thus, we have
$\sum_{e\in E(B^{\prime},W^{\prime})}{\mathrm{charge}(e)}\geq\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(B^{\prime},W^{\prime})\geq\operatorname{dem}_{P}(B,W)-\operatorname{cap}_{G}(B,W).$
@@
By charging an additional unit to each cut edge, we get $\operatorname{dem}_{P}(B,W)\leq\sum_{e\in E(B^{\prime},W^{\prime})}{\mathrm{charge}(e)}$.
@@
We can bound the total charge on each edge as follows. The load $\alpha$ is $O(1)$ for each cluster, as it starts as $\alpha=1$ and is constant after each application of Theorem 6.3. Therefore, each edge (in the subdivision graph) $e\in E(B^{\prime},W^{\prime})$, is charged by Theorem 6.3, $O(\log n\cdot f(S))$ for each refinement-phase cluster $S$ that contains it (i.e., $e\in G^{\prime}[S^{\prime}]$). In particular, an edge $e\in E(B^{\prime},W^{\prime})$ is only charged by clusters $S$ which lie on a root-to-leaf path in $T$.
@@
Consider a path in the decomposition hierarchy $T$, $S_{1}=V,\ldots,S_{k}$, where $S_{i+1}$ is a refinement-phase sub-cluster of $S_{i}$ (that is, we ignore the merge-phase sub-cluster between them). Note that for all $1\leq i<k$, when $S_{i+1}$ was created (as a refinement of some sub-cluster of $S_{i}$), we used $\sigma=|S_{i}|$ in Theorem 6.2. So, $f(S_{i+1})=c_{f}\log\left(\frac{|S_{i}|}{|S_{i+1}|}\right)\cdot\log\log n$. Therefore, it holds that
@@
Therefore, the total charge is $O(\log^{2}n\cdot\log\log n)$ per edge, which gives $\operatorname{dem}_{P}(B,W)=O(\log^{2}n\cdot\log\log n)\cdot\operatorname{cap}_{G^{\prime}}(B^{\prime},W^{\prime})=O(\log^{2}n\cdot\log\log n)\cdot\operatorname{cap}_{G}(B,W)$, as we wanted. Theorem 2.8 then follows from Corollary 5.3. ∎

### v2

@@ 6 Improved Tree Cut-Sparsifier
We begin by re-introducing the relevant notation from Section 5. As per Corollary 5.3, we fix a cut $(U\subseteq V,W=V\setminus U)$ in $G$ and a valid demand state $P$ on $V$, which is $1$-respected by $T$. We work in the subdivision graph $G^{\prime}=(V\cup X_{E},E^{\prime})$. We define for a cluster $S\subseteq V$, the corresponding cluster in the subdivision graph as $S^{\prime}\mathrel{\mathop{:}}=S\cup\{x_{e}\mid e\cap S\geq 1\}$ and let $G^{\prime}[S^{\prime}]$ be the subdivision graph restricted to the set $S^{\prime}$. We let $(U^{\prime}=U\cup\{x_{e}\mid e\cap U\geq 1\},W^{\prime}=V^{\prime}\setminus U^{\prime})$ be the cut corresponding to $(U,W)$ in $G^{\prime}$. We will charge the edges $E(U^{\prime},W^{\prime})$ in $G^{\prime}$ to cover the demand $\operatorname{dem}_{P^{\prime}}(U^{\prime},W^{\prime})$, where $P^{\prime}$ will be the appropriate demand state on $G^{\prime}$.
Algorithm 1 Construction of the tree Cut-Sparsifier
1:function BuildTree($G=(V,E)$)
2:  Create a tree $T$ with root labeled by the cluster $V$.
3:  AltMergeRefine($G,V$)
4:  return $T$
5:end function
6:
7:function AltMergeRefine($G,S$)
8:  if $|S|\leq 1$ then
9:   return
10:  end if
11:  $\sigma\leftarrow|S|$ $\triangleright$ Previous refinement phase cluster, or the root $V$
12:  $(L,R,\{L_{1},\ldots,L_{\ell}\},\{R_{1},\ldots,R_{r}\})\leftarrow\textsc{MergePhase}(G,S,\tau=1)$.
13:  Make $L$ and $R$ children of $S$ in $T$.
14:  for all $C\in\{L_{1},\ldots,L_{\ell}\}\cup\{R_{1},\ldots,R_{r}\}$ do
15:   Make $C$ a child of its corresponding parent ($L$ or $R$).
16:   $\{C_{1},\ldots,C_{k}\}\leftarrow\textsc{Refine}(G,C,\sigma)$ $\triangleright$ partition of $C$
17:   for all $C_{i}\in\{C_{1},\ldots,C_{k}\}$ do
18:     Make $C_{i}$ a child of $C$.
19:     AltMergeRefine($G,C_{i}$).
20:   end for
21:  end for
22:end function
Like in Section 5, we convert the valid demand state $P$ to demand states $P_{\{v\}^{\prime}}$ for each vertex $v\in V$, by distributing the demand vector of $v$ to the split nodes neighboring $v$ evenly. These demand states satisfy Invariant 5.4 with $\alpha=1$. If we denote the sum of these initial demand states as $P_{\mathcal{Q}_{0}}$, Claim 5.8 shows that $\operatorname{dem}_{P}(U,W)\leq\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(U^{\prime},W^{\prime})+\operatorname{cap}_{G}(U,W)$.
@@
Let $P^{(1)}_{S^{\prime}_{i,j}}\mathrel{\mathop{:}}=P_{S^{\prime}_{i,j}}$. We apply the first guarantee of Theorem 6.2 on each of the clusters $S_{i,j}$ ($1\leq i\leq t,1\leq j\leq t_{i}$), to distribute the demand states $P^{(1)}_{S^{\prime}_{i,j}}$ uniformly on the boundary edges of the cluster. Formally, define the average demand vector $N_{i,j}\mathrel{\mathop{:}}=\frac{1}{|X_{B_{i,j}}|}\sum_{v\in X_{B_{i,j}}}{P^{(1)}_{S^{\prime}_{i,j}}(v,\cdot)}$, where $B_{i,j}=E(S_{i,j},V\setminus S_{i,j})$ are the boundary edges of the cluster $S_{i,j}$. Consider the demand state $P^{(2)}_{S^{\prime}_{i,j}}$ which is supported on $X_{B_{i,j}}$, where each demand vector is $N_{i,j}$. By the construction of $P^{(2)}_{S^{\prime}_{i,j}}$, it satisfies Invariant 5.4 for $S_{i,j}$ with the same parameter $\alpha$ as $P^{(1)}_{S^{\prime}_{i,j}}$. In fact, $P^{(2)}_{S^{\prime}_{i,j}}$ is obtained by updating $P^{(1)}_{S^{\prime}_{i,j}}$ by the all-to-all demand matrix on $X_{B_{i,j}}$ scaled by at most $\alpha$ (see Lemma 3.20). Using the first guarantee of Theorem 6.2, and by Facts 3.19(2) and 2.7, the valid demand state $P^{(1)}_{S^{\prime}_{i,j}}-P^{(2)}_{S^{\prime}_{i,j}}$ is $\Omega\left(\frac{1}{\alpha\cdot f(S_{i,j})\cdot\log n}\right)$-respected by $G^{\prime}[S^{\prime}_{i,j}]$. Therefore, by charging $O(\alpha\cdot f(S_{i,j})\cdot\log n)$ to each edge of $E(U^{\prime},W^{\prime})\cap G^{\prime}[S_{i,j}^{\prime}]$, we are able to cover the demand $\operatorname{dem}_{\left(P^{(1)}_{S^{\prime}_{i,j}}-P^{(2)}_{S^{\prime}_{i,j}}\right)}(U^{\prime},W^{\prime})$.
@@
The update by the scaled all-to-all flow problem does not change the total demand (i.e., $P^{(1)}_{S^{\prime}_{i,j}}-P^{(2)}_{S^{\prime}_{i,j}}$ is a valid demand state), so we get
@@
Using Claim 6.4, we have that the total load $P^{(2)}_{S^{\prime}_{i,j}}$ puts on $X_{B_{i,j}}$ is at most $\operatorname{cap}_{G}(B_{i,j})$, and because the demand vectors are the same, we get that the load on each vertex is at most $1$.
@@
Next, we use the second guarantee of Theorem 6.2. This allows us to route in $G^{\prime}[S^{\prime}_{i}]$, one unit of mass from each $x_{f}\in X_{F_{i}}$ to $X_{B_{i}}$, such that each vertex $x_{b}\in X_{B_{i}}$ receives at most $O(1)$ units of mass. Scaling this flow by $2$ gives that we can route two units from each inter-cluster edge to the boundary. This flow can be routed with congestion $O(\log n)$. Hence, by Fact 2.7, we can replace $P^{(2)}_{S^{\prime}_{i}}$ with the updated demand state obtained after this routing, denoted $P^{(3)}_{S^{\prime}_{i}}$, while incurring a charge of $O(\log n)$ on the edges $E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}_{i}]$. Consequently, $P^{(3)}_{S^{\prime}_{i}}$ satisfies Invariant 5.4 for $S_{i}$ with $\alpha=O(1)$, as the demand state is supported on nodes of $X_{B_{i}}$ and the load on each node is constant.
Denote $P^{(3)}_{S^{\prime}}\mathrel{\mathop{:}}=\sum_{i=1}^{t}{P^{(3)}_{S^{\prime}_{i}}}$. We now use Theorem 6.1 on the merge phase partition $\{S_{1},\ldots,S_{t}\}$. It allows us to replace $P^{(3)}_{S^{\prime}}$ with the updated $P^{(4)}_{S^{\prime}}$ which satisfies Invariant 5.4 for $S$ with $\alpha=O(1)$, while charging each edge of $E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$ at most $O(\log n)$ to cover the difference in demands.
Thus, $P^{\mathrm{after}}_{S^{\prime}}\mathrel{\mathop{:}}=P^{(4)}_{S^{\prime}}$ is the demand state required by Theorem 6.3, and the total charge on each edge of $E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$ is dominated by $O(\alpha\cdot f(S_{i,j})\cdot\log n)$. ∎
Using Theorem 6.3, Theorem 2.8 follows exactly like Theorem 2.9 in Section 5. We give the full details for completeness.
@@
Denote by $\mathcal{Q}$ the set of current “active” clusters, such that initially $\mathcal{Q}=\mathcal{Q}_{0}\mathrel{\mathop{:}}=\{\{v\}\mid v\in V\}$. Then in each iteration we remove $\{S_{i,j}\mid 1\leq i\leq t,1\leq j\leq t_{i}\}$ from $\mathcal{Q}$ and add $S$ to $\mathcal{Q}$, where $S_{i,j}$ are the refinement phase sub-clusters as in Theorem 6.3. Let $P_{\mathcal{Q}}$ be the sum of all demand states $P_{S^{\prime}}$, for $S\in\mathcal{Q}$. Using Claim 5.7, we have that $P_{\mathcal{Q}}$ is always a valid demand state on $V^{\prime}$. Claim 5.8 shows that $\operatorname{dem}_{P}(U,W)\leq\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(U^{\prime},W^{\prime})+\operatorname{cap}_{G}(U,W)$.
@@
Consider a replacement of $\mathcal{Q}$ by $\tilde{\mathcal{Q}}=\mathcal{Q}\cup\{S\}\setminus\{S_{i,j}\mid 1\leq i\leq t,1\leq j\leq t_{i}\}$ when we apply Theorem 6.3. Recall that $P^{\mathrm{before}}_{S^{\prime}}=\sum_{i=1}^{t}\sum_{j=1}^{t_{i}}P_{S^{\prime}_{i,j}}$ is replaced by $P^{\mathrm{after}}_{S^{\prime}}$. Let $P^{\mathrm{before}}_{\tilde{\mathcal{Q}}}\mathrel{\mathop{:}}=P_{\mathcal{Q}}$ be the sum of active demand states before the change and $P^{\mathrm{after}}_{\tilde{\mathcal{Q}}}\mathrel{\mathop{:}}=P_{\mathcal{Q}}-P^{\mathrm{before}}_{S^{\prime}}+P^{\mathrm{after}}_{S^{\prime}}$ be the sum after the change. We have $\operatorname{dem}_{\left(P^{\mathrm{before}}_{\tilde{\mathcal{Q}}}-P^{\mathrm{after}}_{\tilde{\mathcal{Q}}}\right)}(U^{\prime},W^{\prime})=\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(U^{\prime},W^{\prime})$. This means that after accumulating all the charges in the applications of Theorem 6.3 on all clusters $S\in T$, the total charge satisfies
$\sum_{e\in E(U^{\prime},W^{\prime})}{\operatorname{charge}(e)}$ $=\sum_{e\in E(U^{\prime},W^{\prime})}\sum_{S\in T}{\operatorname{charge}_{S}(e)}=\sum_{S\in T}\sum_{e\in E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]}{\operatorname{charge}_{S}(e)}$
@@
$\geq\sum_{S\in T}{\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(U^{\prime},W^{\prime})}=\sum_{S\in T}{\operatorname{dem}_{\left(P^{\mathrm{before}}_{\tilde{\mathcal{Q}}}-P^{\mathrm{after}}_{\tilde{\mathcal{Q}}}\right)}(U^{\prime},W^{\prime})}$
$\geq\operatorname{dem}_{\left(P_{\mathcal{Q}_{0}}-P_{V^{\prime}}\right)}(U^{\prime},W^{\prime}).$
@@
where $\operatorname{charge}_{S}(e)$ is the charge incurred on $e$ when applying Theorem 6.3 on the cluster $S$, $\operatorname{charge}(e)$ is the total charge incurred on $e$, and the last inequality uses Fact 3.16.
Since $P_{V^{\prime}}$ satisfies Invariant 5.4(1) for $V$, it follows that $P_{V^{\prime}}=0$. Thus, we have
$\sum_{e\in E(U^{\prime},W^{\prime})}{\operatorname{charge}(e)}\geq\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(U^{\prime},W^{\prime})\geq\operatorname{dem}_{P}(U,W)-\operatorname{cap}_{G}(U,W),$
@@
where the second inequality follows from Claim 5.8. By charging an additional unit to each cut edge, we get $\operatorname{dem}_{P}(U,W)\leq\sum_{e\in E(U^{\prime},W^{\prime})}{\operatorname{charge}(e)}$.
@@
It remains to bound the total charge $\operatorname{charge}(e)$ on each edge $e\in E(U^{\prime},W^{\prime})$. The load $\alpha$ is $O(1)$ for each cluster, as it starts as $\alpha=1$ and is constant after each application of Theorem 6.3. Therefore, each edge (in the subdivision graph) $e\in E(U^{\prime},W^{\prime})$, is charged by Theorem 6.3, $O(\log n\cdot f(S))$ for each refinement phase [^20] Note that the charging in Theorem 6.3 only charges refinement phase clusters. of a cluster $S$ that contains it (i.e., $e\in G^{\prime}[S^{\prime}]$). In particular, an edge $e\in E(U^{\prime},W^{\prime})$ is only charged by clusters $S$ which lie on a root-to-leaf path in $T$.
@@
Consider a path in the decomposition hierarchy $T$, $S_{1}=V,\ldots,S_{k}$, where $S_{i+1}$ is a refinement phase sub-cluster of $S_{i}$ (that is, we consider only refinement phase sub-cluster between them since these are the only charged sub-clusters in Theorem 6.3). Note that for all $1\leq i<k$, when $S_{i+1}$ was created (as a refinement of some sub-cluster of $S_{i}$), we used $\sigma=|S_{i}|$ in Theorem 6.2. So, $f(S_{i+1})=c_{f}\log\left(\frac{|S_{i}|}{|S_{i+1}|}\right)\cdot\log\log n$. Therefore, it holds that
@@
Therefore, the total charge is $O(\log^{2}n\cdot\log\log n)$ per edge, which gives $\operatorname{dem}_{P}(U,W)=O(\log^{2}n\cdot\log\log n)\cdot\operatorname{cap}_{G^{\prime}}(U^{\prime},W^{\prime})=O(\log^{2}n\cdot\log\log n)\cdot\operatorname{cap}_{G}(U,W)$, as we wanted. Theorem 2.8 then follows from Corollary 5.3. ∎